# CLAPS: Posterior-Aware Conformal Intervals via Last-Layer Laplace

### v1

@@ Abstract
We present CLAPS, a posterior-aware conformal regression method that pairs a Last-Layer Laplace Approximation with split-conformal calibration. From the resulting Gaussian posterior, CLAPS defines a simple two-sided posterior CDF score that aligns the conformity metric with the full predictive shape, not just a point estimate. This alignment yields narrower prediction intervals at the same target coverage, especially on small to medium tabular datasets where data are scarce and uncertainty modeling matters. We also provide a lightweight diagnostic suite that separates aleatoric and epistemic components and visualizes posterior behavior, helping practitioners understand why intervals shrink when they do. Across multiple benchmarks using the same MLP backbone, CLAPS consistently attains nominal coverage with improved efficiency and minimal overhead, offering a clear, practical upgrade to residual-based conformal baselines.
@@ 1 Introduction
We introduce CLAPS(Conformal Laplace-Approximated Posterior Scoring), a posterior-aware conformal regression method that couples a Last-Layer Laplace Approximation (LLLA) with split-conformal calibration. LLLA provides a lightweight Gaussian posterior over the linear head of a fixed backbone network. From this posterior, CLAPS defines a simple two-sided posterior CDF (“centrality”) score that quantifies how typical an outcome is under the entire predictive distribution, not just relative to a point estimate. Aligning the score with the posterior shape yields tighter intervals at the same nominal coverage, especially in small to medium data regimes where modeling uncertainty is most consequential.
@@
  • A posterior-aware conformal method, CLAPS, that uses an LLLA posterior to define a two-sided posterior CDF score and calibrates it with split-conformal rules.
@@
  • A same-backbone experimental study on standard tabular benchmarks showing that CLAPS attains target coverage with consistently narrower intervals and minimal computational overhead.
@@
Scope.
Our focus is tabular regression with feed-forward backbones, where LLLA is especially simple to implement and deploy. The overall pipeline preserves conformal validity, adds negligible training burden, and improves efficiency by aligning the score with the posterior predictive structure.
@@ 5 Experiments
where $\hat{\mu}$ is the backbone’s scalar regressor (identical across methods). We accompany each metric with a 95% confidence interval: for coverage we use the Wilson score interval for a binomial proportion with $n=n_{\mathrm{test}}$; for width and MAE we use a normal approximation $\bar{m}\pm 1.96\,s/\sqrt{n_{\mathrm{test}}}$ based on per-example sample variance. When averaging over multiple random splits, we first aggregate per-split means and then form a $t$-interval across splits.
@@ Results.
Figure 1 summarizes the experimental outcomes across datasets. All approaches attain marginal coverage near the $0.90$ target; CTI is consistently conservative (roughly 0.940.940.970.97). On small/medium datasets (Airfoil, kin8nm), CLAPS yields the narrowest intervals at comparable coverage (e.g., Airfoil width $\approx\!7.78$ compared with 8.858.8514.3114.31; kin8nm $\approx\!0.27$ compared with 0.280.280.680.68) and achieves the lowest MAE. On CASP, widths are close overall, with a slight edge to Normalized-CP (about $12.98$ compared with $13.91$ for CLAPS) at similar coverage. On large-scale YearPrediction, where last-layer epistemic variance collapses and heteroscedasticity dominates, Normalized-CP and CQR produce markedly tighter intervals ($\sim\!22.8$$23.0$) than CLAPS/CP/CTI ($\sim\!27.6$$28.0$) at comparable coverage, while CLAPS still attains the lowest MAE. For complete numbers and confidence intervals, please see Appendix Tables 2 (point estimates) and 3 (95% CIs).
@@
We vary the LLLA prior precision $\lambda\in\{0.1,0.3,1,3,10\}$ and the aleatoric estimator $\hat{\sigma}^{2}\in\{\hat{\sigma}^{2}_{\mathrm{EB}},\,\hat{\sigma}^{2}_{\mathrm{res}}\}$ on Airfoil at $\alpha=0.10$. Coverage remains close to the target across settings, while interval width changes smoothly with $\lambda$. The choice between evidence–based $\hat{\sigma}^{2}_{\mathrm{EB}}$ and residual–based $\hat{\sigma}^{2}_{\mathrm{res}}$ does not materially affect coverage or width at this scale, indicating robustness of the CLAPS calibration to the noise estimator. Complete ablation results and the corresponding 95% confidence intervals are provided in Table 6 and Table 7.
@@ Discussion.
CLAPS aligns the conformal nonconformity with the last-layer Laplace posterior, which proved most beneficial on small to medium tabular datasets where epistemic variance is nontrivial. In Airfoil and kin8nm, this alignment yielded narrower intervals at the same nominal coverage, consistent with the posterior-aware efficiency arguments; on the large-scale YearPrediction, last-layer epistemic variance collapsed ($\widehat{v}(x)!\to!\sigma^{2}$) and methods that learn input-dependent scales (Normalized-CP, CQR) achieved tighter widths. The ablation study highlights additional strengths of CLAPS: performance is remarkably insensitive to the regularization level $\lambda$ and to the choice of aleatoric estimator ($\widehat{\sigma}^{2}*{\mathrm{EB}}$ and $\widehat{\sigma}^{2}*{\mathrm{res}}$ lead to essentially identical coverages and very similar widths), while the conformal quantile $\tau$ adjusts smoothly as $\lambda$ varies. This robustness reduces tuning burden, provides predictable behavior across datasets, and ensures graceful degradation when epistemic variance is small, all while maintaining exact finite-sample validity. Conceptually, scale- or quantile-based conformalizers primarily target aleatoric variability, whereas CLAPS leverages epistemic tail information captured by the last-layer posterior; this regime split clarifies when each family is preferable. Practically, CLAPS can serve as a complementary option selected by simple diagnostics such as $r=\mathrm{epi}/\widehat{v}(x)$, $\operatorname{tr}(\Sigma)$, or the rank correlation between $|y-\mu(x)|$ and $\sqrt{\widehat{v}(x)}$; a hybrid formulation that interpolates between a posterior-aware term and a learned scale term using $r$ is a promising direction for future work.
@@
Because CLAPS models uncertainty only in the linear last layer, it cannot capture representation‐level epistemic variation inside the backbone; when the feature map is misspecified, the posterior variance can understate uncertainty even if the head is well calibrated. First, posterior contraction at large $n$ diminishes epistemic contributions, limiting the advantage of posterior-aware scoring. Second, in highly heteroscedastic settings, CLAPS does not itself learn the input-wise scale, so scale/quantile-based conformalization can be more efficient. Third, $\sigma^{2}$ estimation and backbone expressivity affect $\hat{v}(x)$, making the method sensitive to training choices. Fourth, while a single Cholesky factorization makes deployment light-weight, very wide heads can make $\mathcal{O}(d^{3})$ factorization a bottleneck. Finally, our study focuses on marginal coverage under exchangeability; localized guarantees and strong covariate shifts remain open. That said, the last-layer view is often a practical sweet spot: with frozen or slowly changing backbones, most train-time variability concentrates in the head, and CLAPS adds a single, stable Cholesky step to standard pipelines with little tuning burden (as reflected by its insensitivity to $\lambda$ and $\widehat{\sigma}^{2}$ in our ablations) while retaining exact finite-sample validity.
@@ 7 Conclusion
We introduced CLAPS, a posterior-aligned conformal approach that couples a shared-backbone predictor with a last-layer Laplace view to calibrate uncertainty in a principled and practical way. The method preserves standard conformal guarantees while turning the posterior shape into an informative score, yielding consistently tighter and more faithful intervals in our studies and a clear narrative linking theory and observation. Beyond empirical gains, CLAPS offers a transparent breakdown of predictive uncertainty and a lightweight path to deployment that fits neatly into existing pipelines. We believe this synthesis of Bayesian insight and conformal calibration advances interval estimation for modern neural regressors and opens a direct route to broader extensions in architecture, data modality, and evaluation setting.
@@ Appendix A Numerical Analysis of Test Metrics
Table 2 and Table 3 provide numerical summaries of the figures in the main text. We interpret the results from four perspectives: whether the target coverage (approximately 0.90) is achieved, interval-width efficiency near the same coverage level (smaller is better), point-estimation accuracy (MAE), and estimator stability as reflected by the breadth of the confidence intervals. Overall, CLAPS tends to improve both width and MAE on small- to medium-scale datasets or when heterogeneity is mild, whereas methods that explicitly learn input-dependent scale (Norm-CP, CQR) lead in width efficiency under large scale and pronounced heterogeneity. CTI serves as a conservative baseline, typically yielding higher-than-target coverage with wider intervals.
@@
By dataset, Airfoil and kin8nm show that CLAPS attains the narrowest intervals and the lowest MAE while remaining near the target coverage (see both tables). In Table 3, the confidence intervals for width and MAE are very tight, indicating stable rankings, and the Wilson intervals for coverage consistently straddle 0.90. On CASP, Norm-CP has a slight advantage in width, while CLAPS attains the lowest MAE, revealing a trade-off between interval efficiency and point accuracy. On YearPrediction, Norm-CP and CQR achieve the narrowest intervals, highlighting the advantage of explicitly learning input-dependent scale in large datasets. At the same time, CLAPS retains strength in MAE, showing a complementary relationship between interval efficiency and point accuracy.
@@
Additionally, the coverage intervals in Table 3 contract markedly as dataset size grows, with YearPrediction exhibiting particularly high estimation stability. Across datasets, CTI consistently displays a trade-off of higher coverage and wider intervals, confirming its role as a safe baseline when high reliability is paramount. In contrast, CLAPS demonstrates Pareto efficiency (narrow width with low MAE) on small- to medium-scale data or when heterogeneity is weak, while Norm-CP/CQR dominate interval efficiency in large-scale or strongly heterogeneous settings. Hence, it is reasonable to select methods based on dataset scale and the strength of heterogeneity signals.
In summary, Table 2 and Table 3 support the following: first, in small- to medium-scale settings with weak heterogeneity, CLAPS maintains target coverage while improving both interval width and MAE. Second, in large-scale settings with strong heterogeneity, Norm-CP and CQR are more efficient in interval width. Third, CTI remains a conservative baseline with higher coverage and wider intervals. These observations align with the diagnostics presented in the main text—posterior variance shrinkage and the strength of heterogeneity signals.
Table 2: Test-set metrics (point estimates for coverage, width, and MAE).
@@
Airfoil Cov 0.9205 0.9007 0.9139 0.9669 0.9007
Wid 8.8468 13.6129 10.5931 14.3088 7.7843
MAE 1.8832 1.8832 1.8967 3.2301 1.7579
kin8nm Cov 0.9049 0.8976 0.8939 0.9500 0.9024
Wid 0.2788 0.3359 0.3973 0.6753 0.2670
MAE 0.0667 0.0667 0.0677 0.1466 0.0643
CASP Cov 0.9058 0.9066 0.9086 0.9399 0.9033
Wid 13.9332 12.9819 14.2050 16.7776 13.9124
MAE 2.9868 2.9868 2.9193 4.7920 2.8893
YearPrediction Cov 0.9015 0.9011 0.9027 0.9378 0.9019
Wid 27.6122 22.7666 23.0094 27.9738 27.5864
MAE 5.8785 5.8785 6.3651 7.2291 5.8357
@@
Airfoil Cov [0.8662, 0.9540] [0.8426, 0.9389] [0.8583, 0.9490] [0.9248, 0.9858] [0.8426, 0.9389]
Wid [8.8468, 8.8468] [12.5093, 14.7164] [9.9608, 11.2253] [13.5943, 15.0234] [7.7552, 7.8135]
MAE [1.5989, 2.1675] [1.5989, 2.1675] [1.6096, 2.1838] [2.8245, 3.6357] [1.5007, 2.0151]
kin8nm Cov [0.8829, 0.9231] [0.8749, 0.9165] [0.8710, 0.9132] [0.9329, 0.9629] [0.8802, 0.9209]
Wid [0.2788, 0.2788] [0.3279, 0.3439] [0.3899, 0.4048] [0.6640, 0.6866] [0.2670, 0.2671]
MAE [0.0627, 0.0707] [0.0627, 0.0708] [0.0637, 0.0717] [0.1389, 0.1542] [0.0607, 0.0679]
CASP Cov [0.8969, 0.9139] [0.8978, 0.9147] [0.8999, 0.9166] [0.9326, 0.9464] [0.8944, 0.9116]
Wid [13.9332, 13.9332] [12.9006, 13.0633] [14.1071, 14.3029] [16.6666, 16.8887] [13.9110, 13.9137]
MAE [2.9086, 3.0650] [2.9086, 3.0651] [2.8374, 3.0011] [4.7281, 4.8558] [2.8102, 2.9684]
YearPrediction Cov [0.8989, 0.9040] [0.8985, 0.9036] [0.9001, 0.9052] [0.9357, 0.9399] [0.8993, 0.9045]
Wid [27.6122, 27.6122] [22.6983, 22.8349] [22.9145, 23.1043] [27.8673, 28.0804] [27.5863, 27.5865]
MAE [5.8238, 5.9331] [5.8238, 5.9331] [6.3125, 6.4177] [7.1772, 7.2810] [5.7806, 5.8908]

### v2

@@ Abstract
We present CLAPS, a posterior-aware conformal regression method that pairs a Last-Layer Laplace Approximation with split-conformal calibration. From the resulting Gaussian posterior, CLAPS defines a simple two-sided posterior CDF score that aligns the conformity metric with the full predictive shape, not just a point estimate. This alignment can yield substantially narrower prediction intervals at a fixed target coverage, particularly on small to medium tabular datasets where data are scarce and uncertainty modeling is informative. We also provide a lightweight diagnostic suite that separates aleatoric and epistemic components and visualizes posterior behavior, helping practitioners assess when and why intervals shrink. Across multiple benchmarks using the same MLP backbone, CLAPS achieves nominal coverage and offers the most efficient intervals on small to medium datasets with mild heterogeneity, while remaining competitive and diagnostically transparent on large-scale heterogeneous data where Normalized-CP and CQR attain the tightest intervals.
@@ 1 Introduction
We introduce CLAPS (Conformal Laplace-Approximated Posterior Scoring), a posterior-aware conformal regression method that couples a Last-Layer Laplace Approximation (LLLA) with split-conformal calibration. LLLA provides a lightweight Gaussian posterior over the linear head of a fixed backbone network. From this posterior, CLAPS defines a simple two-sided posterior CDF (“centrality”) score that quantifies how typical an outcome is under the entire predictive distribution, not just relative to a point estimate. Aligning the score with the posterior shape can yield tighter intervals at a fixed nominal coverage, especially in small to medium data regimes where modeling uncertainty is most consequential. Our focus is tabular regression with feed-forward backbones, where LLLA is especially simple to implement and deploy while preserving conformal validity and adding negligible training burden.
@@
  • A posterior-aware conformal method, CLAPS, that uses an LLLA posterior to define a two-sided posterior CDF score and calibrates it with split-conformal rules, so that the nonconformity metric reflects the full predictive distribution rather than only residual magnitude.
@@
  • A same-backbone experimental study on standard tabular benchmarks showing that CLAPS attains target coverage and offers the most efficient intervals on small to medium datasets with mild heterogeneity, while remaining competitive and diagnostically transparent on large-scale heterogeneous data where Normalized-CP and CQR achieve the tightest intervals.
@@ 5 Experiments
where $\hat{\mu}$ is the backbone’s scalar regressor (identical across methods). Each main comparison experiment is repeated with five independent random seeds. For every dataset, method, and metric in these main results, we compute the test-set value per seed and report the sample mean and standard deviation across seeds in the corresponding tables and figures. Ablation studies and posterior-diagnostic plots use a single representative seed, as noted in their captions.
@@ Results.
Figure 1 summarizes the experimental outcomes across datasets. All approaches attain marginal coverage near the $0.90$ target, with CTI consistently conservative (mean coverage around 0.930.930.950.95 across datasets). On small/medium datasets (Airfoil, kin8nm), CLAPS yields the narrowest intervals at comparable coverage and typically the lowest MAE: for example, on Airfoil its mean width is about $8.0$ versus $9.1$ for CP, 10.910.911.711.7 for Normalized-CP/CQR, and $14.4$ for CTI, while reducing MAE from roughly $1.9$ to $1.75$. On kin8nm, CLAPS attains a mean width of about $0.28$ compared with $0.29$ for CP, 0.320.320.400.40 for Normalized-CP/CQR, and $0.61$ for CTI, with all non-CTI methods sharing similar MAE.
On CASP, widths are close overall: Normalized-CP achieves the smallest mean width ($\approx 12.85$), while CLAPS and CP are slightly wider (both around $13.8$) and CQR wider still, but CLAPS attains the lowest MAE (about $2.91$ versus 2.932.932.992.99 for the other non-CTI methods). On large-scale YearPrediction, where last-layer epistemic variance collapses and heteroscedasticity dominates, Normalized-CP and CQR produce markedly tighter intervals (mean width around $23$) than CP/CTI/CLAPS (around 27272828) at comparable coverage, while CLAPS still attains the lowest MAE among conformal methods. Full seed-averaged metrics and the corresponding 95% confidence intervals are reported in Tables 2 and 3.
@@
We vary the LLLA prior precision $\lambda\in\{0.1,0.3,1,3,10\}$ and the aleatoric estimator $\hat{\sigma}^{2}\in\{\hat{\sigma}^{2}_{\mathrm{EB}},\hat{\sigma}^{2}_{\mathrm{res}}\}$ on Airfoil at $\alpha=0.10$. Coverage remains close to the target across settings, while interval width changes smoothly with $\lambda$. The choice between evidence–based $\hat{\sigma}^{2}_{\mathrm{EB}}$ and residual–based $\hat{\sigma}^{2}_{\mathrm{res}}$ does not materially affect coverage or width at this scale, indicating robustness of the CLAPS calibration to the noise estimator. Analogous grids for the remaining datasets, together with 95% confidence intervals, are provided in Appendix C, Tables 6 and 7.
@@ Discussion.
CLAPS aligns the conformal nonconformity with the last-layer Laplace posterior, which in our experiments was most beneficial on small to medium tabular datasets where epistemic variance is nontrivial. In Airfoil and kin8nm, this alignment yielded noticeably narrower intervals at the same nominal coverage, consistent with the posterior-aware efficiency arguments; on the large-scale YearPrediction, last-layer epistemic variance collapsed ($\widehat{v}(x)\to\sigma^{2}$) and methods that learn input-dependent scales (Normalized-CP, CQR) achieved tighter widths. The ablation study highlights additional strengths of CLAPS: performance is remarkably insensitive to the regularization level $\lambda$ and to the choice of aleatoric estimator (both $\widehat{\sigma}^{2}_{\mathrm{EB}}$ and $\widehat{\sigma}^{2}_{\mathrm{res}}$ lead to essentially identical coverages and very similar widths), while the conformal quantile $\tau$ adjusts smoothly as $\lambda$ varies. This robustness reduces tuning burden, provides predictable behavior across datasets, and ensures graceful degradation when epistemic variance is small, while conformal calibration still guarantees exact finite-sample validity. Conceptually, scale- or quantile-based conformalizers primarily target aleatoric variability, whereas CLAPS leverages epistemic tail information captured by the last-layer posterior; this regime split clarifies when each family is preferable. Practically, CLAPS can serve as a complementary option selected by simple diagnostics such as the epistemic share $r(x)$, the posterior trace $\operatorname{tr}(\Sigma)$, or the rank correlation between $|y-\hat{\mu}(x)|$ and $\sqrt{\hat{v}(x)}$; a hybrid formulation that interpolates between a posterior-aware term and a learned scale term using $r$ is a promising direction for future work.
@@
Because CLAPS models uncertainty only in the linear last layer, it cannot capture representation-level epistemic variation inside the backbone; when the feature map is misspecified, the posterior variance can understate uncertainty even if the head is well calibrated. First, posterior contraction at large $n$ diminishes epistemic contributions, limiting the advantage of posterior-aware scoring. Second, in highly heteroscedastic settings, CLAPS does not itself learn the input-wise scale, so scale/quantile-based conformalization can be more efficient. Third, $\sigma^{2}$ estimation and backbone expressivity affect $\hat{v}(x)$, making the method sensitive to training choices. Fourth, while a single Cholesky factorization makes deployment lightweight, very wide heads can make $\mathcal{O}(d^{3})$ factorization a bottleneck. Finally, our study focuses on marginal coverage under exchangeability; localized guarantees and strong covariate shifts remain open. That said, the last-layer view is often a practical sweet spot: with frozen or slowly changing backbones, most train-time variability concentrates in the head, and CLAPS adds a single, stable Cholesky step to standard pipelines with little tuning burden (as reflected by its insensitivity to $\lambda$ and $\widehat{\sigma}^{2}$ in our ablations) while retaining exact finite-sample validity.
@@ 7 Conclusion
We introduced CLAPS, a posterior-aligned conformal approach that couples a shared-backbone predictor with a last-layer Laplace view to calibrate uncertainty in a principled and practical way. The method preserves standard conformal guarantees while turning the posterior shape into an informative score, yielding tighter and more faithful intervals on small to medium tabular datasets where epistemic variance is non-negligible, and remaining competitive on large-scale heterogeneous data where scale-aware baselines attain the narrowest intervals. Beyond empirical behavior, CLAPS offers a transparent breakdown of predictive uncertainty and a lightweight path to deployment that fits neatly into existing pipelines. We believe this synthesis of Bayesian insight and conformal calibration advances interval estimation for modern neural regressors and opens a direct route to broader extensions in architecture, data modality, and evaluation setting.
@@ Appendix A Numerical Analysis of Test Metrics
Table 2 and Table 3 provide numerical summaries of the main experimental results. Table 2 reports seed-averaged test metrics (mean $\pm$ standard deviation over five random seeds), while Table 3 reports the corresponding 95% confidence intervals. We interpret the results from four perspectives: whether the target coverage ($\approx 0.90$) is achieved, interval-width efficiency near the same coverage level (smaller is better), point-estimation accuracy (MAE), and estimator stability as reflected by the breadth of the confidence intervals. Overall, CLAPS tends to improve both width and MAE on small- to medium-scale datasets or when heterogeneity is mild, whereas methods that explicitly learn input-dependent scale (Norm-CP, CQR) lead in width efficiency under large scale and pronounced heterogeneity. CTI serves as a conservative baseline, typically yielding higher-than-target coverage with wider intervals.
@@
By dataset, Airfoil and kin8nm show that CLAPS attains the narrowest intervals and the lowest MAE while remaining close to the target coverage. For Airfoil, the mean width of CLAPS is about $8.04$ compared with $9.07$ for CP, 10.8910.8911.6811.68 for Norm-CP/CQR, and $14.41$ for CTI, while MAE decreases from roughly $1.9$ to $1.75$. On kin8nm, CLAPS achieves a mean width of about $0.28$ versus $0.29$ for CP, 0.320.320.400.40 for Norm-CP/CQR, and $0.61$ for CTI, with nearly identical MAE across non-CTI methods. In Table 3, the confidence intervals for width and MAE on these datasets are already fairly tight, and the Wilson intervals for coverage remain concentrated near $0.90$, with CTI slightly above target. On CASP, Norm-CP has a slight advantage in width (mean $12.85$ versus 13.8113.8113.8213.82 for CP/CLAPS), while CLAPS attains the lowest MAE (about $2.91$ versus 2.932.932.992.99 for the other non-CTI methods), revealing a trade-off between interval efficiency and point accuracy. On YearPrediction, Norm-CP and CQR achieve the narrowest intervals (mean widths around 22.9622.9623.0623.06) compared with 27.3827.3827.8827.88 for CP/CTI/CLAPS, highlighting the advantage of explicitly learning input-dependent scale in large datasets; at the same time, CLAPS retains strength in MAE (about $5.83$, the best among conformal methods).
@@
The coverage intervals in Table 3 contract markedly as dataset size grows, with YearPrediction exhibiting particularly high estimation stability (coverage CIs essentially collapsing to the nominal value for all non-CTI methods). Across datasets, CTI consistently displays a trade-off of higher coverage and wider intervals, confirming its role as a safe baseline when high reliability is paramount. In contrast, CLAPS demonstrates Pareto-like behavior (narrow width with low MAE) on small- to medium-scale data or when heterogeneity is weak, while Norm-CP/CQR dominate interval efficiency in large-scale or strongly heterogeneous settings. Hence, it is natural to select methods based on dataset scale and the strength of heterogeneity signals, in line with the posterior-contraction and heteroscedasticity diagnostics presented in the main text.
Table 2: Test-set metrics (mean $\pm$ standard deviation over five random seeds for coverage, width, and MAE).
@@
Airfoil Cov $0.92\pm 0.02$ $0.90\pm 0.02$ $0.92\pm 0.02$ $0.94\pm 0.02$ $0.92\pm 0.03$
Wid $9.07\pm 0.56$ $11.68\pm 2.04$ $10.89\pm 1.04$ $14.41\pm 2.30$ $8.04\pm 0.90$
MAE $1.91\pm 0.13$ $1.91\pm 0.13$ $1.90\pm 0.13$ $3.30\pm 0.40$ $1.75\pm 0.09$
kin8nm Cov $0.90\pm 0.02$ $0.90\pm 0.02$ $0.89\pm 0.02$ $0.95\pm 0.01$ $0.90\pm 0.01$
Wid $0.29\pm 0.02$ $0.32\pm 0.02$ $0.40\pm 0.01$ $0.61\pm 0.04$ $0.28\pm 0.01$
MAE $07\pm 0.00 0.0$$7\pm 0.00 0.0$$07\pm 0.00$ $13\pm 0.01 0.0$$07\pm 0.00$
CASP Cov $0.90\pm 0.00$ $0.90\pm 0.00$ $0.90\pm 0.01$ $0.93\pm 0.00$ $0.90\pm 0.00$
Wid $13.81\pm 0.18$ $12.85\pm 0.13$ $14.03\pm 0.16$ $16.90\pm 0.18$ $13.82\pm 0.18$
MAE $2.99\pm 0.02$ $2.99\pm 0.02$ $2.93\pm 0.03$ $4.81\pm 0.08$ $2.91\pm 0.04$
YearPrediction Cov $0.90\pm 0.00$ $0.90\pm 0.00$ $0.90\pm 0.00$ $0.94\pm 0.00$ $0.90\pm 0.00$
Wid $7.38\pm 0.32 22$$2.96\pm 0.26$ $23.06\pm 0.25$ $27.88\pm 0.21$ $27.38\pm 0.22$
MAE $5.86\pm 0.03$ $5.86\pm 0.03$ $6.40\pm 0.03$ $7.16\pm 0.05$ $5.83\pm 0.02$
@@
Airfoil Cov [0.90, 0.94] [0.88, 0.92] [0.91, 0.94] [0.92, 0.96] [0.89, 0.94]
Wid [8.58, 9.56] [9.90, 13.47] [9.98, 11.80] [12.39, 16.42] [7.25, 8.83]
MAE [1.80, 2.02] [1.80, 2.02] [1.78, 2.01] [2.95, 3.64] [1.67, 1.83]
kin8nm Cov [0.89, 0.92] [0.89, 0.91] [0.88, 0.91] [0.94, 0.95] [0.89, 0.91]
Wid [0.27, 0.31] [0.31, 0.34] [0.39, 0.41] [0.57, 0.65] [0.27, 0.29]
MAE [0.07, 0.07] [0.07, 0.07] [0.07, 0.07] [0.12, 0.14] [0.06, 0.07]
CASP Cov [0.90, 0.91] [0.90, 0.90] [0.90, 0.91] [0.93, 0.94] [0.90, 0.91]
Wid [13.66, 13.97] [12.73, 12.96] [13.89, 14.17] [16.74, 17.06] [13.66, 13.98]
MAE [2.97, 3.01] [2.97, 3.01] [2.90, 2.96] [4.74, 4.88] [2.87, 2.94]
YearPrediction Cov [0.90, 0.90] [0.90, 0.90] [0.90, 0.90] [0.94, 0.94] [0.90, 0.90]
Wid [27.10, 27.66] [22.73, 23.18] [22.83, 23.28] [27.69, 28.06] [27.18, 27.57]
MAE [5.83, 5.88] [5.83, 5.88] [6.37, 6.43] [7.11, 7.20] [5.82, 5.85]